Solve the recurrence hm12– 2h9+1+hn=hı (0) + 2 (n > 0), with initial values ho = 1 and = 1.
Answer #1

solutions characteristic equation. 7?_97 + 20 a bravo اراء (( tm = cont cancion -, tcan for non- homogeneous part, As right s* (n+2)tB (megy? + c (n+2]? +D9n+2 – 24194104_28 (m«13e - 2((n+15-9 Dont & Ant+Britan² +Dqh mu ņ + A[metan? taun 16) +B[r3 +128 = (roeff, of m?) 24A +GB 2-1/2 (co-eff an IHA +6B+26=0 (co-eft of constant) D=1 Coefficient of 2n) | 39 63=\q=1 5o B=33/2ishi- citat 1 - 2 + 1 ta 120 tat 1-6tilth8 24 1 = 22 +54 - C2=154 SO -. tm = fantaun Intl a tan CScanned with CamScanner